SA12. Illustrate the path of a projectile launched at a velocity of 150.0 m/s at an angle of 50O with the horizontal and find, illustrate, and label the following in the same illustration you have. a. The viy of the projectile b. The vix of the projectile c. Its vy 2.0 seconds after it was launched d. Its vf based on letter c e. Its maximum height (𝐦𝐚𝐱 𝐝𝐲 ↑) f. The time it will reach a height of 3.0 m (𝒕 ↑) g. Its position/location 4.0 seconds of its flight (𝐝𝐲 ↑) h. Its range ( R )


1
Expert's answer
2021-12-06T09:46:55-0500

a)

"v_y=v\sin \alpha=114.9~\frac ms,"

b)

"v_x=v\cos\alpha=96.4~\frac ms,"

c)

"v_{y2}=v_y-gt=95.3~m,"

d)

"v_{y0}=0~\frac ms,"

e)

"h=\frac{v_y^2}{2g}=667.5~m,"

f)

"s=v_yt-\frac{gt^2}2,\implies"

"t=0.03~s" or "t=23.42~s,"

g)

"h_4=v_yt-\frac{gt^2}2=381.2~m,"

h)

"R=v_xt_0=226.6~m."